the amount of money you will receive from the bank now, if the bank approved the loan with redemption of r30 000 due 8 months from now with an applicable amount with sample discount rate of 16.5 per year is?


1
Expert's answer
2022-02-06T14:22:59-0500

"PV=\frac{FV}{(1+i)^n}"


n=8


"i=\frac{0.16}{12}=0.0133"


"PV=\frac{30000}{(1+0.0133)^8}=26990.87"